# Assessing the dynamical assumptions in Tsirelson inequality tests of non-classicality in harmonic oscillators

### v2

@@ Appendix A Tsirelson Quantity from the Wigner Function
Thus, The Tsirelson quantity can also be written in terms of the Wigner quasiprobability distribution $W_{\rho}(x,p)$ as
$\langle A\rangle=\text{Tr}(\rho A)=\int\int dx\ dp\ W_{A}(x,p)W_{\rho}(x,p)\ ,$ (A1)
where $W_{A}(x,p)$ (and likewise $W_{\rho}(x,p)$) is the phase-space representation of the operator $A$, obtained from its Weyl transform:
$W_{A}(x,p)=\int_{-\infty}^{\infty}dy\,e^{-ipy/\hbar}\left\langle x+\frac{y}{2}\middle|A\middle|x-\frac{y}{2}\right\rangle\ .$ (A2)
@@ Appendix B Single crossing dynamical assumption
An alternative set of inequalities can be formulated under the single crossing assumption (SCA), which is an alternative dynamical assumption that instead asserts that p(+,,+)=p(,+,)=0p(+,-,+)=p(-,+,-)=0, forbidding that the particle cross the origin more than once. A very rapidly oscillating system could cross many times in between measurements whilst maintaining p(+,,+)=p(,+,)=0p(+,-,+)=p(-,+,-)=0, but we assume that the relative amplitudes of these erratic trajectories are small.
SCA can be imposed in the same way as was done uniform precession in the original Tsirelson derivation, by setting p(+,,+)=p(,+,)=0p(+,-,+)=p(-,+,-)=0. Using the moment expansions 1.7, it is straightforward to obtain
@@
Classically, one of the simplest examples of a system for which the SCA is valid would be a free particle. Acted upon by no external forces, a free particle will only ever cross a pre-defined origin at most once. In the case of a uniformly-precessing system, such as the harmonic oscillator, this assumption holds true over half a period. We choose measurement times $t_{i}=\frac{iT}{6}$, $i=0,1,2$, where $T$ is the time period of the oscillator.
@@
Since $\langle A\rangle=\langle\hat{Q}_{t=0}\rangle+\langle\hat{Q}_{t=1/3}\rangle+\langle\hat{Q}_{t=2/3}\rangle$ and $\langle A_{\text{SCA}}\rangle=\langle\hat{Q}_{t=0}\rangle-\langle\hat{Q}_{t=1/6}\rangle+\langle\hat{Q}_{t=1/3}\rangle$, we see that these two quantities only differ by one term. If indeed it is true that $\langle A\rangle=\langle A_{\text{SCA}}\rangle$ then it is implied that $\langle\hat{Q}_{t=2/3}\rangle=-\langle\hat{Q}_{t=1/6}\rangle$. In fact, the more general result holds true that $\langle\hat{Q}_{t}\rangle=-\langle\hat{Q}_{t+1/2}\rangle\ \forall\ t$. That is to say, the single-time average of the dichotomous variable $Q$ merely flips sign after half a period.
We can understand why this is the case by considering what happens to the energy eigenfunctions after exactly half a period. Undergoing unitary time evolution in accordance with the QHO Hamiltonian, the eigenstate $|n\rangle$ accrues a relative phase of $e^{-i\pi n}$ after half a period. As such, the odd-$n$ eigenstates will pick up a minus sign, whilst the even-$n$ states are unaffected. In the language of wavefunctions, it is simple to see that this means a reflection across the $x$-axis after exactly half a period. Therefore, quantum mechanically, the quantities $\langle A\rangle$ and $\langle A_{\text{SCA}}\rangle$ are identical.
@@ Appendix C Matrix Elements in an Energy Eigenbasis
$|\Psi\rangle=\sum_{n=0}^{\infty}c_{n}|n\rangle,\qquad\hat{Q}=2\hat{\theta}-\mathds{1},$ (C1)
where $\hat{\theta}$ is the Heaviside projector on $x>0$. To obtain explicit matrices we truncate the Hilbert space at a finite excitation $N$:
@@
$Q_{nk}=\!\int_{-\infty}^{\infty}\!\!\Psi_{n}(x)\Psi_{k}(x)\,\bigl[2\theta(x)-1\bigr]\,\mathrm{d}x\ ,$ (C4)
using the property that the energy eigenstates for the 1D harmonic oscillator can always be taken to be real to drop the complex conjugate on $\Psi_{n}(x)$. To make the integral tractable, we split the integral at the origin and use the overlap formula in the appendix of Ref.[65],
@@
where $F_{nk}=\langle n|\hat{\theta}|k\rangle$ and $F_{nn}=\frac{1}{2}$. Therefore, we find
@@
To evaluate Equation C5, we use the closed form expressions for the eigenstates of the harmonic oscillator in terms of the Hermite polynomials $H_{n}(x)$,
@@
Finally, since $\hat{U}|n\rangle=e^{-2\pi in/3}|n\rangle$,
@@
We verified this implementation by checking matrix elements of the $F_{nk}$ agreed with those cited in Ref.[65] and our implementation of the Tsirelson matrix was verified by checking some known violating states.
@@ Appendix D Diagonalisation of the Tsirelson operator in an energy eigenbasis
in Eq. 2.8 vanishes unless the difference (nk)(n-k) is a multiple of 3. Moreover, not only is the Zaw–Scarani state the maximally violating state in the $N=6$ subspace, but also in the $N=7$ and $N=8$ subspaces. For $N=9$, on the other hand, we can achieve a superior violation of $\langle A\rangle\approx 1.12$ with the state
@@
Fig. 7 shows the results of a spectral decomposition of the Tsirelson operator in a high-energy truncated subspace. This serves as a good illustration that the classical values $A=\pm 1$ are “accumulation points”, but violations past this value are indeed possible. None of our computed eigenvalues exceed $|A|\approx 1.26$, which agrees with the value of the maximum possible violation quoted in Tsirelson’s original paper [1]. A more accurate estimate of the true upper and lower bounds on the spectral norm of $A$ has been obtained in the recent work of Ref. [9]. Fig. 8 illustrates that considering a progressively larger subspace from the low-lying energy states upwards does indeed lead to an asymptotic convergence towards Tsirelson’s quoted value (up to an arbitrary scaling factor) for the maximum violation in the full QHO Hilbert space.
Figure 7: A histogram of the spectrum of eigenvalues of $A$ is plotted by using an energy truncation of $E=70(\hbar\omega)$. The y-axis refers to the number of eigenvalues found within a given range.
Figure 8: The computed maximal eigenvalue of A in an increasingly large subspace, tending towards a finite maximal violation. Note, importantly, that we re-scale the eigenvalues as $\frac{1}{6}(A+3)$. This is only an arbitrary scaling choice to see agreement with the definition of the operator considered by Tsirelson in Ref. [2]. The Zaw–Scarani state ($|\Psi_{Z}\rangle$) is pointed out at $N=6$.
We have also computed the eigenvectors of $A$ in increasingly larger subspaces, catalogued in Table 3, and have found that they all obey the same pattern of being composed solely of energy eigenstates that are multiples of three.
@@
90 $|\Psi\rangle\rightarrow$ True eigenstate of $\hat{A}$ 1.2131
@@ Appendix E Superpositions of Two Energy Eigenstates
As established in Ref.[1], single energy eigenstates of the harmonic oscillator cannot violate Tsirelson’s inequality due to their definite parity. The next logical step is to find violations produced by a superposition of two eigenstates. However, as we show now, any superposition of two energy eigenstates is insufficient to produce a violation.
@@ E.1 Proof Strategy
$\langle A\rangle=|\alpha|^{2}\underbrace{\langle n|\hat{A}|n\rangle}_{0}+|\beta|^{2}\underbrace{\langle k|\hat{A}|k\rangle}_{0}+2\operatorname{Re}(\alpha^{*}\beta)\langle n|\hat{A}|k\rangle.$ (E2)
@@ Appendix F Calculation of correlators in the quantum harmonic oscillator
The measurement operator in the Heisenberg picture is given by $\hat{Q}(t)=e^{iHt}(2\hat{\theta}-1)e^{-iHt}$, where we choose $t_{2}-t_{1}=\tau$ and $t_{1}=0$, such that
$\hat{Q}_{j}=e^{iH(j-1)\tau}(2\hat{\theta}-1)e^{-iH(j-1)\tau}\ .$ (F2)
@@
$\therefore C_{12}=\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|\hat{Q}_{1}e^{i\hat{H}\tau}\hat{Q}_{1}e^{-i\hat{H}\tau}|k\rangle+\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|e^{i\hat{H}\tau}\hat{Q}_{1}e^{-i\hat{H}\tau}\hat{Q}_{1}|k\rangle$ (F4)
@@
$e^{i\hat{H}\tau}=\sum_{m=0}^{\infty}e^{im\tau}|m\rangle\langle m|\ ,$ (F5)
@@ Appendix G All Three Correlators are Equal on $\mathcal{H}_{k}$
by leveraging the unitarity of $\hat{U}$ and its action on $|\psi_{k}\rangle$. We then group some terms in the aim of showing $F_{12k}=F_{23k}=F_{13k}$,
$F_{ijk}=\operatorname{Re}\omega^{k(j-i)}\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{i-j}\hat{Q}_{1}|\psi_{k}\rangle\ .$ (G3)
Clearly, $F_{12k}=\operatorname{Re}\omega^{k}\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-1}\hat{Q}_{1}|\psi_{k}\rangle=F_{23k}$ but it also quickly follows that $F_{12k}=F_{13k}$. Note,
$F_{13k}=\operatorname{Re}\omega^{2k}\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-2}\hat{Q}_{1}|\psi_{k}\rangle\ ,$ (G4)
where $\omega^{2k}=(\omega^{k})^{*}$, $\hat{U}^{-2}=\hat{U}$ and $\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-2}\hat{Q}_{1}|\psi_{k}\rangle=\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-1}\hat{Q}_{1}|\psi_{k}\rangle^{*}$. So,
$F_{13k}=\operatorname{Re}\left(\omega^{k}\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-1}\hat{Q}_{1}|\psi_{k}\rangle\right)^{*}=\operatorname{Re}\omega^{k}\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-1}\hat{Q}_{1}|\psi_{k}\rangle=F_{12k}=F_{23k}\ .$ (G5)
@@ Appendix H Calculation of sequential measurement probabilities
and define the Heisenberg-picture projection operators $P_{+}(t)=\hat{\theta}(t)$ and $P_{-}(t)=(1-\hat{\theta}(t))$, where we have defined the time-evolved Heaviside projector onto the positive x-axis:
$\hat{\theta}(t)\equiv\int_{0}^{\infty}e^{-iHt}|x\rangle\langle x|e^{iHt}\mathrm{d}x\ .$ (H2)
@@
$p_{123}(+,+,+)=\mathrm{Tr}(e^{2iH\tau}P_{+}(t_{1})e^{-iH\tau}P_{+}(t_{1})e^{-iH\tau}P_{+}(t_{1})\rho P_{+}(t_{1})e^{iH\tau}P_{+}(t_{1})e^{-iH\tau})\ .$ (H4)
Then, writing
@@
We can alternatively consider making the first measurement a “weak measurement”, which leads to the expression
$p_{(1)23}(+,+,+)=\frac{1}{2}\text{Tr}\left(P_{+}(t_{3})P_{+}(t_{2})(P_{+}(t_{1})\rho+\rho P_{+}(t_{1}))P_{+}(t_{2})\right)=\frac{1}{2}\text{Tr}\left(P_{+}(t_{3})P_{+}(t_{2})\left\{P_{+}(t_{1}),\rho\right\}P_{+}(t_{2})\right)\ .$ (H9)
Classically, this distribution is identical to $p_{123}(+,+,+)$; but quantum mechanically, the sum rule (which is not generally valid) does indeed hold over the weak first measurement:
$\sum_{s_{1}}p_{23}(s_{1},s_{2},s_{3})=p_{23}(s_{2},s_{3})\ .$ (H10)
This distribution therefore could be said to lie between the intrusive three-time measurement probability, $p_{123}$, and the quasi-probability distribution, $q$ (where the latter is equivalent to making all three measurements weak). In this case,
$p_{(1)23}(+,+,+)-p_{(1)23}(-,-,-)\equiv\Delta p_{(1)23}=\frac{1}{4}(\langle A\rangle+D+I)\ ,$ (H11)
where the interference terms would only stem from the second measurement: $I=\langle Q_{3}^{(2)}\rangle-\langle Q_{3}\rangle$. We could just as well consider a distribution with a weak second measurement, where
$p_{1(2)3}(+,+,+)=\frac{1}{2}\text{Tr}\left(P^{(3)}_{+}\{P^{(2)}_{+},P^{(1)}_{+}\rho P^{(1)}_{+}\}\right)\ ,$ (H12)
and so
$p_{1(2)3}(+,+,+)=\frac{1}{2}\text{Tr}\left(e^{2iH\tau}\hat{\theta}e^{-2iH\tau}\{e^{iH\tau}\hat{\theta}e^{-iH\tau},\hat{\theta}\rho\hat{\theta}\}\right)\\ =\frac{1}{2}\text{Tr}\left(e^{2iH\tau}\hat{\theta}e^{-iH\tau}\hat{\theta}e^{iH\tau}\hat{\theta}\rho\hat{\theta}\right)+\frac{1}{2}\text{Tr}\left(e^{2iH\tau}\hat{\theta}e^{-2iH\tau}\hat{\theta}\rho\hat{\theta}e^{iH\tau}\hat{\theta}e^{-iH\tau}\right)\ ,$ (H13)
@@
$\frac{1}{2}\text{Tr}\left(e^{2iH\tau}\hat{\theta}e^{-iH\tau}\hat{\theta}e^{iH\tau}\hat{\theta}\rho\hat{\theta}\right)=\frac{1}{32}\sum_{n,k,p,q,r}c_{n}^{*}c_{k}e^{i\tau(2p+r-q)}L_{pq}L_{qm}L_{mk}L_{np}\ .$ (H14)

### v3

@@ Appendix A Tsirelson Quantity from the Wigner Function
The Tsirelson quantity can also be written in terms of the Wigner quasiprobability distribution $W_{\rho}(x,p)$ as
$\langle A\rangle=\text{Tr}(\rho\hat{A})=\int\int dx\ dp\ W_{A}(x,p)W_{\rho}(x,p)\ ,$ (A1)
where $W_{A}(x,p)$ (and likewise $W_{\rho}(x,p)$) is the phase-space representation of the operator $\hat{A}$, obtained from its Weyl transform:
$W_{A}(x,p)=\int_{-\infty}^{\infty}dy\,e^{-ipy/\hbar}\left\langle x+\frac{y}{2}\middle|\hat{A}\middle|x-\frac{y}{2}\right\rangle\ .$ (A2)
@@ Appendix B Single crossing dynamical assumption
An alternative set of inequalities can be formulated under the single crossing assumption (SCA), which is an alternative dynamical assumption that instead asserts that p(+,,+)=p(,+,)=0p(+,-,+)=p(-,+,-)=0, forbidding that the particle cross the origin more than once. A very rapidly oscillating system could cross many times in between measurements whilst maintaining p(+,,+)=p(,+,)=0p(+,-,+)=p(-,+,-)=0, but we assume that the relative amplitudes of these erratic trajectories are small. In fact, similar inequalities can be obtained using the polytope method along with dynamical constraints over a full period as shown in Ref. [plavala_tsirelson_2024].
SCA can be imposed in the same way as was done uniform precession in the original Tsirelson derivation, by setting p(+,,+)=p(,+,)=0p(+,-,+)=p(-,+,-)=0. Using the moment expansions 2.3, it is straightforward to obtain
@@
Classically, one of the simplest examples of a system for which the SCA is valid would be a free particle. Acted upon by no external forces, a free particle will only ever cross a pre-defined origin at most once. In the case of a uniformly-precessing system, such as the harmonic oscillator, this assumption holds true over half a period. We choose measurement times $t_{j}=\frac{jT}{6}$, $j=0,1,2$, where $T$ is the time period of the oscillator.
@@
Since $\langle A\rangle=\langle{Q}_{t=0}\rangle+\langle{Q}_{t=1/3}\rangle+\langle{Q}_{t=2/3}\rangle$ and $\langle A_{\text{SCA}}\rangle=\langle{Q}_{t=0}\rangle-\langle{Q}_{t=1/6}\rangle+\langle{Q}_{t=1/3}\rangle$, we see that these two quantities only differ by one term. If indeed it is true that $\langle A\rangle=\langle A_{\text{SCA}}\rangle$ then it is implied that $\langle{Q}_{t=2/3}\rangle=-\langle{Q}_{t=1/6}\rangle$. In fact, the more general result holds true that $\langle{Q}_{t}\rangle=-\langle{Q}_{t+1/2}\rangle\ \forall\ t$. That is to say, the single-time average of the dichotomic variable $Q$ merely flips sign after half a period.
We can understand why this is the case by considering what happens to the energy eigenfunctions after exactly half a period. Undergoing unitary time evolution in accordance with the QHO Hamiltonian, the eigenstate $|n\rangle$ accrues a relative phase of $e^{-i\pi n}$ after half a period. As such, the odd-$n$ eigenstates will pick up a minus sign, whilst the even-$n$ states are unaffected. In the language of wavefunctions, it is simple to see that this means a reflection across the $y$-axis after exactly half a period. Therefore, quantum mechanically, the quantities $\langle A\rangle$ and $\langle A_{\text{SCA}}\rangle$ are identical.
@@ Appendix C Matrix Elements in an Energy Eigenbasis
$|\Psi\rangle=\sum_{n=0}^{\infty}c_{n}|n\rangle,\qquad\hat{Q}=2\theta(\hat{x})-\mathds{1},$ (C1)
[…]
@@
$Q_{nk}=\!\int_{-\infty}^{\infty}\!\!\Psi_{n}(x)\Psi_{k}(x)\,\bigl[2\theta(\hat{x})-1\bigr]\,\mathrm{d}x\ ,$ (C4)
using the property that the energy eigenstates for the 1D harmonic oscillator can always be taken to be real to drop the complex conjugate on $\Psi_{n}(x)$. To make the integral tractable, we split the integral at the origin and use the overlap formula in the appendix of Ref.[MawbySHO],
@@
where $F_{nk}=\langle n|\theta(\hat{x})|k\rangle$ and $F_{nn}=\frac{1}{2}$. Therefore, we find
@@
To evaluate Eq. C5, we use the closed form expressions for the eigenstates of the harmonic oscillator in terms of the Hermite polynomials $H_{n}(x)$,
@@
Finally, since $U|n\rangle=e^{-2\pi in/3}|n\rangle$,
@@
We verified this implementation by checking matrix elements of the $F_{nk}$ agreed with those cited in Ref.[MawbySHO] and our implementation of the Tsirelson matrix was verified by checking some known violating states.
@@ Appendix D Diagonalisation of the Tsirelson operator in an energy eigenbasis
in Eq. 3.7 vanishes unless the difference (nk)(n-k) is a multiple of 3. Moreover, not only does the Zaw–Scarani state produce the largest violations state in the $N=6$ subspace, but also in the $N=7$ and $N=8$ subspaces. For $N=9$, on the other hand, we can achieve a superior violation of $\langle A\rangle\approx 1.12$ with the state
@@
Fig. 7 shows the results of a spectral decomposition of the Tsirelson operator in a high-energy truncated subspace. This serves as a good illustration that the classical values $A=\pm 1$ are “accumulation points” [tsirelson_how_2006], but violations past this value are indeed possible. None of our computed eigenvalues exceed $\approx 1.26$, which agrees with the value of the maximum possible violation quoted in Tsirelson’s original paper [tsirelson_how_2006]. A more accurate estimate of the true upper and lower bounds on the spectral norm of $\hat{A}$ has been obtained in the recent work of Ref. [zaw2024threeanglevariantstsirelsonsprecession]. Fig. 8 illustrates that considering a progressively larger subspace from the low-lying energy states upwards does indeed lead to an asymptotic convergence towards Tsirelson’s quoted value (up to an arbitrary scaling factor) for the maximum violation in the full QHO Hilbert space.
Figure 7: A histogram of the spectrum of eigenvalues of $\hat{A}$ is plotted by using an energy truncation of $E=70(\hbar\omega)$. The y-axis refers to the number of eigenvalues found within a given range.
Figure 8: The computed maximal eigenvalue of $\hat{A}$ in an increasingly large subspace, tending towards a finite maximal violation. Note, importantly, that we re-scale the eigenvalues $\lambda$ as $\frac{1}{6}(\lambda+3)$. This is only an arbitrary scaling choice to see agreement with the definition of the operator considered by Tsirelson in Ref. [plavala_tsirelson_2024]. The Zaw–Scarani state ($|\Psi_{Z}\rangle$) is pointed out at $N=6$.
We have also computed the eigenvectors of $\hat{A}$ in increasingly larger subspaces, catalogued in Table 3, and have found that they all obey the same pattern of being composed solely of energy eigenstates that are multiples of three.
@@
90 $|\Psi\rangle\rightarrow$ True eigenstate of A{A} 1.2131
@@ Appendix E Superpositions of Two Energy Eigenstates
As established in Ref.[tsirelson_how_2006], single energy eigenstates of the harmonic oscillator cannot violate Tsirelson’s inequality due to their definite parity. The next logical step is to find violations produced by a superposition of two eigenstates. However, as we show now, any superposition of two energy eigenstates is insufficient to produce a violation.
@@ E.1 Proof Strategy
$\langle A\rangle=|\alpha|^{2}\underbrace{\langle n|A|n\rangle}_{0}+|\beta|^{2}\underbrace{\langle k|A|k\rangle}_{0}+2\operatorname{Re}(\alpha^{*}\beta)\langle n|A|k\rangle.$ (E2)
@@ Appendix F Calculation of correlators in the quantum harmonic oscillator
The measurement operator in the Heisenberg picture is given by $\hat{Q}(t)=e^{iHt}(2{\theta(\hat{x})}-1)e^{-iHt}$, where we choose $t_{2}-t_{1}=\tau$ and $t_{1}=0$, such that
$\hat{Q}_{j}=e^{iH(j-1)\tau}(2{\theta}(\hat{x})-1)e^{-iH(j-1)\tau}\ .$ (F2)
@@
$\therefore C_{12}=\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|\hat{Q}_{1}e^{i{H}\tau}\hat{Q}_{1}e^{-i{H}\tau}|k\rangle+\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|e^{i{H}\tau}\hat{Q}_{1}e^{-i{H}\tau}\hat{Q}_{1}|k\rangle$ (F4)
@@
$e^{i{H}\tau}=\sum_{m=0}^{\infty}e^{im\tau}|m\rangle\langle m|\ ,$ (F5)
@@ Appendix G All Three Correlators are Equal on $\mathcal{H}_{k}$
by leveraging the unitarity of $U$ and its action on $|\psi_{k}\rangle$. We then group some terms in the aim of showing $F_{12k}=F_{23k}=F_{13k}$,
$F_{ijk}=\operatorname{Re}\omega^{k(j-i)}\langle\psi_{k}|\hat{Q}_{1}U^{i-j}\hat{Q}_{1}|\psi_{k}\rangle\ .$ (G3)
Clearly, $F_{12k}=\operatorname{Re}\omega^{k}\langle\psi_{k}|\hat{Q}_{1}U^{-1}\hat{Q}_{1}|\psi_{k}\rangle=F_{23k}$ but it also quickly follows that $F_{12k}=F_{13k}$. Note,
$F_{13k}=\operatorname{Re}\omega^{2k}\langle\psi_{k}|\hat{Q}_{1}U^{-2}\hat{Q}_{1}|\psi_{k}\rangle\ ,$ (G4)
where $\omega^{2k}=(\omega^{k})^{*}$, $U^{-2}=U$ and $\langle\psi_{k}|\hat{Q}_{1}U^{-2}\hat{Q}_{1}|\psi_{k}\rangle=\langle\psi_{k}|\hat{Q}_{1}U^{-1}\hat{Q}_{1}|\psi_{k}\rangle^{*}$. So,
$F_{13k}=\operatorname{Re}\left(\omega^{k}\langle\psi_{k}|\hat{Q}_{1}U^{-1}\hat{Q}_{1}|\psi_{k}\rangle\right)^{*}=\operatorname{Re}\omega^{k}\langle\psi_{k}|\hat{Q}_{1}U^{-1}\hat{Q}_{1}|\psi_{k}\rangle=F_{12k}=F_{23k}\ .$ (G5)
@@ Appendix H Calculation of sequential measurement probabilities
and define the Heisenberg-picture projection operators $P_{+}(t)={\theta}(\hat{x}(t))$ and $P_{-}(t)=[1-{\theta}(\hat{x}(t))]$, in order to make more concise the notation for the time-evolved Heaviside projector onto the positive x-axis:
$P_{+}(t)\equiv{\theta}(\hat{x}(t))=\int_{0}^{\infty}e^{-iHt}|x\rangle\langle x|e^{iHt}\mathrm{d}x\ .$ (H2)
@@
$p_{123}(+,+,+)=\mathrm{Tr}(e^{2iH\tau}P_{+}(t_{1})e^{-iH\tau}P_{+}(t_{1})e^{-iH\tau}P_{+}(t_{1})\rho P_{+}(t_{1})e^{iH\tau}P_{+}(t_{1})e^{-iH\tau})\ ,$ (H4)
recognising that $P_{+}(t_{1})$ is simply the Heaviside operator without any time evolution. is Then, writing
@@
There are also two other quasi-probability distributions, as elucidated in the discussion surrounding Eq. 2.8. We construct these by beginning from Eq. 2.5 and either replacing $C_{13}$ by $C_{13}^{(2)}$ (the first possibility) or replacing $C_{23}$ by $C_{23}^{(1)}$ (the second). (Doing both would effectively yield Eq. 2.7.) Thus, we define
$q_{A}(+,+,+)=\frac{1}{2}\text{Tr}\left(P_{+}(t_{3})P_{+}(t_{2})\left\{P_{+}(t_{1}),\rho\right\}P_{+}(t_{2})\right)\ .$ (H9)
Classically, this distribution is identical to $p_{123}(+,+,+)$. In this case,
$q_{A}(+,+,+)-q_{A}(-,-,-)\equiv\Delta q_{A}=\frac{1}{4}(\langle A\rangle+D+I)\ ,$ (H10)
where the interference terms would only stem from the second measurement: $I=\langle Q_{3}^{(2)}\rangle-\langle Q_{3}\rangle$. Secondly,
$q_{B}(+,+,+)=\frac{1}{2}\text{Tr}\left(P_{+}(t_{3})\{P_{+}(t_{2}),P_{+}(t_{1})\rho P_{+}(t_{1})\}\right)\ .$ (H11)
For reference, this is readily computed in terms of unevolved Heaviside operators:
[…]
@@
$\frac{1}{2}\text{Tr}\left(e^{2iH\tau}{\theta}(\hat{x})e^{-iH\tau}{\theta}(\hat{x})e^{iH\tau}{\theta}(\hat{x})\rho{\theta}(\hat{x})\right)=\frac{1}{32}\sum_{n,k,p,q,r}c_{n}^{*}c_{k}e^{i\tau(2p+r-q)}L_{pq}L_{qm}L_{mk}L_{np}\ .$ (H13)